A population consists of the numbers of 1-5.List all the possible samples of size 3 from this population and construct the sampling distribution of the sample mean.

1
Expert's answer
2022-05-17T08:52:44-0400

We have population values 1,2,3,4,5, population size N=5 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{1+2+3+4+5}{5}=3"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{N}=\dfrac{4+1+0+1+4}{5}=2""\sigma=\sqrt{\sigma^2}=\sqrt{2}\approx1.4142"


The number of possible samples which can be drawn without replacement is "^{N}C_n=^{5}C_3=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 1,2,3 & 6/3 \\
 \hdashline
 2 & 1,2,4 & 7/3 \\
 \hdashline
 3 & 1,2,5 & 8/3\\
 \hdashline
 4 & 1,3,4 & 8/3 \\
 \hdashline
 5 & 1,3,5 & 9/3 \\
 \hdashline
 6 & 1,4,5 & 10/3 \\
 \hdashline
 7 & 2,3,4 & 9/3 \\
 \hdashline
 8 & 2,3,5 & 10/3 \\
 \hdashline
 9 & 2,4,5 & 11/3 \\
 \hdashline
 10 & 3,4,5 & 12/3 \\
 \hdashline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X})& \bar{X}^2f(\bar{X}) 
\\ \hline
6/3 & 1/10 & 6/30 & 36/90 \\
 \hdashline
 7/3 & 1/10 & 7/30 & 49/90 \\
 \hdashline
 8/3 & 2/10 & 16/30 & 128/90 \\
 \hdashline
 9/3 & 2/10 & 18/30 & 162/90 \\
 \hdashline
 10/3 & 2/10 & 20/30 & 200/90 \\
 \hdashline
 11/3 & 1/10 & 11/30 & 121/90 \\
 \hdashline
 12/3 & 1/10 & 12/30 & 144/90 \\
 \hdashline
\end{array}"



Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=3=\mu"



The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{840}{90}-(3)^2=\dfrac{1}{3}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{\dfrac{1}{3}}\approx0.57735"